OpenStudy (anonymous):

Which function table is correct for the rule y=x over 2 + 4

OpenStudy (anonymous):

You can pick from these charts

thomaster (thomaster):

This function?\[y=\frac{ x }{ 2+4 }\]

OpenStudy (anonymous):

No the +4 is on the outside

thomaster (thomaster):

\[y=\frac{ x }{ 2 }+4\]

thomaster (thomaster):

allright fill in 12 for x. what is the answer?

OpenStudy (anonymous):

Yes

thomaster (thomaster):

@heyheyhelpme ?

OpenStudy (anonymous):

8

thomaster (thomaster):

no. 12/2=6 +4 = 10

thomaster (thomaster):

So for answer 1, the first x value is correct, now the next one is 4 what do you get for (4/2)+4?

OpenStudy (anonymous):

I thought you did it all at one time?

thomaster (thomaster):

yes but you need to x/2 first then add the +4

thomaster (thomaster):

so what is y when x=4?

OpenStudy (anonymous):

10

thomaster (thomaster):

how did you come to 10? \[\frac{ 4 }{ 2 }+4 = 2+4 = 6\]

OpenStudy (anonymous):

OHHHHHHHHHH

thomaster (thomaster):

if you're doing this with a calculator, do it like this:\[\left( \frac{ x }{ 2 } \right)+4\]

thomaster (thomaster):

so for x=12, y=10 and for x=4, y=6. for option a there is 1 x value left, 0. What is y when x=0?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

1

thomaster (thomaster):

(0/2)+4 = 1 ??????

OpenStudy (anonymous):

No that's not right

thomaster (thomaster):

how many pieces of apple do you get when you cut NO apple in 2?

OpenStudy (anonymous):

0

thomaster (thomaster):

thats true :) now you add 4 apples to the 0 apples you already have :P

OpenStudy (anonymous):

4

thomaster (thomaster):

yep, what are the y values of funtion table a?

OpenStudy (anonymous):

Table A

thomaster (thomaster):

yes x=12, y=10 x=4, y=6 x=0, y=4 So the answer to you question is function table A is correct for the rule y= (x/2)+4

thomaster (thomaster):

all of the other tables have at least 1 x values that do not have the correct value for y

OpenStudy (anonymous):

KK I have more

OpenStudy (anonymous):

Identify the rule for the function table. y = 3x y = x + 24<This one y = x – 24 y = 2x + 12

thomaster (thomaster):

x=12, y=36 x=20, y=44 x=76, y=100 Lets test the first option: y=3x fill in 12 for x: y=3*12 The answer is 36 so the y value is correct Next one: x=20 --> y=3*20 The answer is 60, so the first option is not the rule of the function table second one: y=x+24 y=12+24=36 y=20+24=44 y=76+24=100 All y values are correct, so y=x+24 is the winning function :)

OpenStudy (anonymous):

Which rule matches the function graphed below? y = x + 3 y = x – 4 y = x – 3 y = x + 4

OpenStudy (anonymous):

+3

thomaster (thomaster):

Do you have a graphical calculator? If not just put the functions in google it will plot them for you. You'll see that Y=X-3 is correct.

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Oh

thomaster (thomaster):

I'd love to help you further but it's 2:30 am so i'm going to sleep now :P

OpenStudy (anonymous):

In the morning

thomaster (thomaster):

yes 2:30 after midnight

OpenStudy (anonymous):

Okay goodnight